OpenStudy (anonymous):

what percent of110 is 99

OpenStudy (dumbcow):

90%

OpenStudy (watchmath):

10 percent of 110 is 11 So 90 percent of 110 is 99

OpenStudy (angela210793):

x/100*110=99 x/100=99/110 x=90

OpenStudy (anonymous):

(99/110)*100 =90%

OpenStudy (anonymous):

see i got 90 but if you divide 9 into 99you get 11

OpenStudy (anonymous):

but i didn't multiply by 100